OpenStudy (anonymous):

Can some one walk me through this equation: 8= (x-3)^3/4

OpenStudy (anonymous):

Step 1: Exponentiate both sides of the equation to the power of 4. This will get rid of the 1/4 power. Step 2: Root both sides to the power of 3. This will get rid of the cube power. Step 3: Add 3 to both sides. This will get simply the variable side. Step 4: By now, you should have your x-value which I assume is what you needed.

OpenStudy (anonymous):

Okay give me a sec.

OpenStudy (anonymous):

not to butt in, but i would take the cubed root first, before i raised to the power of 4 the numbers are much easier that way

OpenStudy (anonymous):

\[8=(x-3)^{\frac{3}{4}}\iff 8^{\frac{4}{3}}=x-3\]

OpenStudy (anonymous):

So I would get x= 5^3/4 right @satellite73

OpenStudy (anonymous):

@Cardinal_Carlo Doing it your way I got \[x=-5\pm8i \sqrt{3}\] Was that right?

OpenStudy (anonymous):

Not quite. My approach was more of a step-by-step breakdown of Satellite's approach. You should've gotten what he got.

OpenStudy (anonymous):

\[8^{\frac{ 4 }{ 3 }} + 3 = x\] See?

OpenStudy (anonymous):

Yeah but there is a bit of a problem and that is that none of the given answers matches that answer. @Cardinal_Carlo

OpenStudy (anonymous):

If that's the case, from now on, please provide an answer list as you post your question. It makes it easier for all of us.

OpenStudy (anonymous):

Oh okay, sorry for that. @Cardinal_Carlo The answers that they provided me are. 16 19 13 11

OpenStudy (anonymous):

\[8^{\frac{ 4 }{ 3 }} + 3 = ~?\] The answer is actually one of the choices you put up.

OpenStudy (anonymous):

Hint: Simplify the expression.